### v1

# dyGRASS: Dynamic Spectral Graph Sparsification via Localized Random Walks on GPUs

Abstract
This work presents dyGRASS, an efficient dynamic algorithm for spectral sparsification of large undirected graphs that undergo streaming edge insertions and deletions. At its core, dyGRASS employs a random-walk-based method to efficiently estimate node-to-node distances in both the original graph (for decremental update) and its sparsifier (for incremental update). For incremental updates, dyGRASS enables the identification of spectrally critical edges among the updates to capture the latest structural changes. For decremental updates, dyGRASS facilitates the recovery of important edges from the original graph back into the sparsifier. To further enhance computational efficiency, dyGRASS employs a GPU-based non-backtracking random walk scheme that allows multiple walkers to operate simultaneously across various target updates. This parallelization significantly improves both the performance and scalability of the proposed dyGRASS framework. Our comprehensive experimental evaluations reveal that dyGRASS achieves approximately a 10$\times$ speedup compared to the state-of-the-art incremental sparsification (inGRASS) algorithm while eliminating the setup overhead and improving solution quality in incremental spectral sparsification tasks. Moreover, dyGRASS delivers high efficiency and superior solution quality for fully dynamic graph sparsification, accommodating both edge insertions and deletions across a diverse range of graph instances originating from integrated circuit simulations, finite element analysis, and social networks.
I Introduction
Graph-based methodologies play a fundamental role in addressing numerous challenges within electronic design automation (EDA), including logic synthesis [1, 2], verification [3], layout optimization [4], static timing analysis (STA) [5], circuit partitioning [6], circuit modeling [7], and simulation [8]. Graph sparsification techniques, in particular, have been shown to be effective in accelerating a variety of compute-intensive EDA tasks, such as circuit simulations [9, 10, 11, 12, 13], allowing more efficient vector-less integrity verification of power grids [14, 15], and identifying worst-case on-chip temperature distributions [16].
Mathematics and theoretical computer science researchers have long studied the problem of simplifying large graphs through spectral graph theory [17, 18, 19, 20]. Recent investigations into spectral graph sparsification have yielded significant advancements, allowing for constructing substantially sparser subgraphs while preserving the key graph spectral properties such as the first few eigenvalues and eigenvectors of the graph Laplacian. These breakthroughs have facilitated the development of nearly linear time algorithms for solving large sparse matrices and partial differential equations (PDEs) [21, 22]. Additionally, they have empowered various applications in graph-based semi-supervised learning (SSL), computing stationary distributions of Markov chains and personalized PageRank vectors, spectral graph partitioning, data clustering, max flow and multi-commodity flow of undirected graphs, and the implementation of nearly-linear time circuit simulation and verification algorithms [23, 18, 11, 24, 16].
While existing spectral graph sparsification methods are scalable for extracting sparsified graphs from scratch [11, 12, 25, 26, 13], they are not suitable for handling cases with streaming edge insertions and deletions efficiently, which may pose a significant bottleneck in the modern EDA workflows. For example, the iterative nature of chip design means that even minor adjustments to the circuit schematic or layout, such as adding or removing circuit components, require the spectral sparsifiers to be updated with each change. In practical applications like power grid design and optimization [27], adding or removing a small number of metal wires requires reconstructing the entire spectral sparsifier, which quickly becomes computationally expensive. While recent theoretical advancements in dynamic graph sparsification algorithms offer promise [28, 29, 30, 31], the practical applicability of these techniques in real-world EDA workflows remains an open question.
Figure 1: The proposed dynamic spectral sparsification method (dyGRASS) and its applications in (non)linear circuit network modeling and optimization tasks.
To address this gap, we propose a practically efficient algorithmic framework, dyGRASS, for dynamic spectral sparsification of large undirected graphs. By leveraging localized random walks, dyGRASS allows updating the spectral sparsifier in constant time for each edge insertion/deletion, as shown in Fig. 1. In each edge update, dyGRASS estimates the spectral distortion [11, 12, 26] due to edge insertion/deletion by leveraging a localized random walk algorithm: for edge insertions dyGRASS determines whether to retain or remove edges bounded by a resistance diameter, whereas for edge deletions dyGRASS recovers edges along a selected path with minimum resistance to retain structural properties in the sparsifier.
TABLE I: Feature Comparison: inGRASS vs. dyGRASS
Aspect inGRASS [32] dyGRASS
Setup Time $O(n\log n)$ N/A
Update Time $O(\log n)$ $O(1)$
Incremental Update
Decremental Update
Fully Dynamic
GPU Acceleration
As shown in Table I where $n$ denotes the number of vertices in the graph, the state-of-the-art inGRASS framework requires $O(n\log n)$ time for the setup phase and supports only incremental edge insertions [32], whereas dyGRASS eliminates the costly setup phase and process each edge insertion or deletion in constant time, offering a more efficient and comprehensive approach to dynamic spectral graph sparsification. Moreover, unlike inGRASS that keeps using the resistance embedding of the initial sparsifier obtained during the setup phase for processing newly added edges, dyGRASS offers a fully dynamic sparsification framework for continuously exploiting the latest graph structure for achieving a nearly-optimal sparsification solution. In addition, our dyGRASS framework allows for utilizing GPUs to handle large-scale updates efficiently, achieving substantial speedups for processing large dynamically evolving graphs. The key contributions of this work have been summarized as follows:
  1. 1.
    We propose dyGRASS for dynamically updating spectral graph sparsifiers. Unlike the latest inGRASS framework [32] that requires an expensive setup phase and supports only a limited number of incremental edge insertions, dyGRASS can handle any size of edge insertions/deletions with greatly improved efficiency.
  2. 2.
    The proposed dyGRASS framework leverages localized random walks to efficiently approximate graph distances and assess spectral distortion caused by each update. Since dyGRASS is a strongly local algorithm, it enjoys a constant time complexity for each edge insertion/deletion operation.
  3. 3.
    By accelerating random walks on Graphics Processing Units (GPUs), dyGRASS achieves a 10×10$\times$10 × runtime speedup over the CPU-based implementation and 2×2$\times$2 × speedup over the state-of-the-art method [32] on real-world large graphs, while preserving solution quality.
II Background
II-A Graph Laplacian, quadratic form, and spectral similarity
For a weighted undirected graph $G=(V,E,w)$, $V$ ($|V|=n$) and $E$ represent the sets of vertices and edges, respectively, and $w$ denotes a positive weight function with $w_{i,j}$ denoting the edge weight. Let $A$ denote the $n\times n$ adjacency matrix, then the Laplacian matrix $L_{G}$ is defined as $L_{G}:=D-A$, where $D$ is the degree matrix. The eigenvalues and eigenvectors of $L_{G}$ offer insights into graph structure properties such as the number of connected components and vertex connectivity. The Laplacian quadratic form $x^{\top}L_{G}x$ is a fundamental tool in spectral graph theory, enabling analysis of graph cuts, clustering, and conductance.
Spectral similarity between the Laplacians of graphs $A$ and $B$ can be quantified using the following inequality [33]:
$\frac{x^{\top}L_{A}x}{\epsilon}\leq x^{\top}L_{B}x\leq\epsilon x^{\top}L_{A}x,$ (1)
where $L_{A}$ and $L_{B}$ represent the Laplacian matrices of $A$ and $B$, with a parameter $\epsilon>0$. A smaller $\epsilon$ or relative condition number $\kappa(L_{A},L_{B})$ indicates greater spectral similarity between $A$ and $B$.
II-B Spectral sparsification
Spectral sparsification methods are used to represent a large graph in a more concise form by removing redundant or less important edges while preserving its essential structural and spectral properties. Given a graph $G=(V,E,w)$ with its Laplacian matrix $L_{G}$, its sparsifier $H=(V,E^{\prime},w^{\prime})$ is its subgraph with its Laplacian matrix $L_{H}$, where $|E^{\prime}|\ll|E|$, while preserving the original spectral graph properties, such as the Laplacian eigenvalues and eigenvectors. Spectral sparsification retains these properties, enabling efficient algorithms and analyses while reducing computational costs. Spectral sparsification has been applied to efficiently solve large systems of equations and tackle graph-based computational problems. For example, instead of directly solving the original Laplacian matrix of graph $G$, the Laplacian of the sparsified graph $H$ can be used as a preconditioner in standard iterative sparse matrix algorithms for reducing total solution time [12].
II-C Effective resistance
Definition 1
The effective resistance between vertices $(p,q)\in V$ is defined as
$R^{eff}_{p,q}=b_{p,q}^{\top}L^{\dagger}b_{p,q}=\sum\limits_{i=2}^{|V|}\frac{(u_{i}^{\top}b_{p,q})^{2}}{u_{i}^{\top}Lu_{i}},$ (2)
where $L^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of the graph Laplacian matrix $L$, $u_{i}\in\mathbb{R}^{|V|}$ for $i=1,...,|V|$ denote the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\lambda_{i}$ for $i=1,...,|V|$, ${b_{p}}\in\mathbb{R}^{|V|}$ denotes the standard basis vector with all zero entries, with the $p$-th entry equal to 1, and ${b_{p,q}}=b_{p}-b_{q}$.
Computing effective resistances directly according to Definition 1 proves impractical for large graphs. However, our dynamic algorithm only requires fast estimations of edge effective resistances, which can be achieved using a highly scalable, constant-time random walk approach.
II-D Random walk
A random walk on a graph is a stochastic process where a walker starts at a vertex and moves to a randomly chosen neighbor at each step. Random walks are a fundamental tool in graph analysis, offering insights into structural and spectral properties and enabling sampling, distance estimation, and connectivity analysis without full graph exploration [34]. Let $G=(V,E)$ be an undirected graph, with $|V|=n$ vertices and $|E|=m$ edges. For any two vertices $p,q\in V$, the hitting time $H_{p,q}$ is the expected number of steps for a walker starting at $p$ to reach $q$ for the first time.
Lemma II.1
For any two vertices $p$ and $q$ in $G$, the commute time $C_{p,q}$ is the expected round-trip time between $p$ and $q$, which is directly proportional to the effective resistance distance between $p$ and $q$: $C_{p,q}=H_{p,q}+H_{q,p}=2m\cdot R^{eff}_{p,q}$ [35].
The above relationship underscores the connection between random walk dynamics and graph spectral properties, which allows for an efficient approximation of effective resistance using random walk metrics.
However, using a Monte Carlo approach to approximate commute times directly remains impractical for large graphs, as the computational effort scales with the total edge count $m$.
Recent work by Peng et al. [36] proposed local algorithms for estimating effective resistance by approximating random walk commute times and Laplacian pseudo-inverses. Their approach, such as EstEff-TranProb, uses a series expansion of the transition matrix and performs multiple $\ell$-step random walks to estimate multi-hop transition probabilities. While theoretically elegant and suitable for graphs with small mixing time, the required number of steps $\ell$ grows quickly as the spectral gap closes. In graphs with high spectral radius, this makes the method impractical for large graphs due to the need for long-range random walks. Moreover, the reliance on long-range walks makes them costly for online or streaming settings. In contrast, our method avoids dependence on global mixing properties or pseudo-inverse approximations. By leveraging localized non-backtracking random walks, we construct efficient upper bounds on effective resistance, enabling fast edge filtering with minimal exploration. This makes our approach more robust in the face of slow mixing and more suitable for dynamic, large-scale graphs.
II-E Dynamic graph algorithms
Dynamic graph algorithms are designed to efficiently handle graphs that undergo updates over time. These algorithms can be categorized as follows:
  1. 1.
    Incremental algorithms: These algorithms efficiently handle additions of edges to the graph. They update the data structure representing the graph to incorporate the new elements while minimizing the computational cost.
  2. 2.
    Decremental algorithms: These algorithms efficiently handle edge deletions from the graph. They update the data structure to reflect the removal of elements while preserving the integrity of the graph.
  3. 3.
    Fully dynamic algorithms: These algorithms handle both additions and deletions of edges from the graph. They dynamically adjust the data structure to accommodate changes while maintaining efficient query performance.
This work focuses on an efficient yet fully dynamic graph update algorithm, leveraging a localized random-walk-based distance estimation technique to manage edge insertions and deletions within the spectral graph sparsifier.
III dyGRASS: Dynamic Graph Sparsification
III-A Overview of dyGRASS
The proposed dyGRASS algorithm dynamically updates the spectral sparsifier in response to streaming edge updates (insertions and deletions), as illustrated in Fig.2. The algorithm takes as input the original graph $G^{(0)}$ and its initial spectral sparsifier $H^{(0)}$, which is precomputed using the GRASS method [12]. Our approach comprises two main phases: (1) incremental spectral graph sparsification and (2) decremental spectral graph sparsification. By integrating these two phases, the dyGRASS algorithm efficiently updates the initial spectral sparsifier $H^{(0)}$ without recomputing the sparsifier from scratch, thereby significantly reducing computational overhead.
Figure 2: Overview of the dyGRASS framework, with dashed lines indicating the graph where localized random walks are initialized.
III-B Incremental spectral sparsification with dyGRASS
(a) Spectral Impact of Edge Insertions: Suppose an edge $e_{p,q}\notin E^{(l)}$ is newly introduced between vertices $p$, $q$ during the $l$-th iteration, updating the original graph $G^{(l)}=(V,E^{(l)},w^{(l)})$. When incorporating $e_{p,q}$ into current sparsifier $H^{(l)}=(V,E^{\prime(l)},w^{\prime(l)})$, the eigenvalue perturbation for each eigenvector of $H^{(l)}$ can be assessed by the following lemma [26].
Lemma III.1
Let $H=(V,E^{\prime},w^{\prime})$, where $w^{\prime}:E^{\prime}\rightarrow\mathbb{R}+$, denote the sparsified weighted graph of $G$, and $L_{H}$ denotes its Laplacian matrix. The $i$-th Laplacian eigenvalue perturbation due to $\delta L_{H}=w^{\prime}_{p,q}b_{pq}b^{\top}_{pq}$ can be computed as:
$\delta\lambda_{i}=w^{\prime}_{p,q}(u_{i}^{\top}b_{pq})^{2},$ (3)
where $u_{i}$ represents the eigenvector corresponding to the $i$-th eigenvalue $\lambda_{i}$ of the Laplacian matrix $L_{H}$.
Thus, if $e_{q,p}$ has a larger $w^{\prime}_{p,q}(u_{i}^{\top}b_{pq})^{2}$, it is considered spectrally critical to $\lambda_{i}$. This indicates that including this edge in the updated sparsifier would cause a larger perturbation to the $i$-th eigenvalue $\lambda_{i}$ and its corresponding eigenvector $u_{i}$.
Lemma III.2
Construct a weighted eigensubspace matrix $U_{K}$ for $K$-dimensional spectral graph embedding using the first $K$ Laplacian eigenvectors and eigenvalues as follows:
$U_{K}=\left[\frac{u_{2}}{\sqrt{\lambda_{2}}},...,\frac{u_{K}}{\sqrt{\lambda_{K}}}\right],$ (4)
then the spectral distortion of the new edge $e_{p,q}$ will become the total $K$-eigenvalue perturbation $\Delta_{K}$ when $K\rightarrow N$[26]:
$\Delta_{K}=\sum\limits_{i=2}^{{K}}\frac{\delta{{\lambda}_{i}}}{\lambda_{i}}=w_{p,q}\|U_{K}^{\top}b_{pq}\|^{2}_{2}\approx w_{p,q}R(p,q).$ (5)
Lemma 3.2 defines the spectral distortion $\Delta_{K}$ of an edge as the product of its edge weight and effective resistance between its two end vertices. We next show that the overall distortion remains bounded by the condition number $\kappa(L_{G},L_{H})$.
Consider the generalized eigenvalue problem for the matrix pencil $(L_{H},L_{G})$ and apply the Courant-Fischer theorem. Specifically, we have:
$\max_{\begin{subarray}{c}p,q\in V\\ p\neq q\end{subarray}}\frac{b_{p,q}^{\top}L_{H}^{+}\,b_{p,q}}{b_{p,q}^{\top}L_{G}^{+}\,b_{p,q}}\;\;\leq\;\;\max_{\begin{subarray}{c}\|v\|\neq 0\\ v^{\top}\mathbf{1}=0\end{subarray}}\frac{v^{\top}L_{H}^{+}\,v}{v^{\top}L_{G}^{+}\,v}\;=\;\lambda_{\max},$ (6)
where $b_{p,q}$ is the incidence vector of edge $e_{p,q}$, and $\mathbf{1}$ is the all-one vector. From
$\frac{b_{p,q}^{\top}L_{H}^{+}\,b_{p,q}}{b_{p,q}^{\top}L_{G}^{+}\,b_{p,q}}\;=\;\frac{R_{p,q,H}}{R_{p,q,G}},$ (7)
we see that taking the maximum over all node pairs $(p,q)$ implies
$\max_{p,q}\,\frac{R_{p,q,H}}{R_{p,q,G}}\;\;\leq\;\;\lambda_{\max}(L_{G}\,L_{H}^{+}).$ (8)
Furthermore, to find $\lambda_{\min}(L_{G},L_{H})$, one applies the Courant-Fischer theorem again:
$1\;\leq\;\lambda_{\min}=\;\min_{\begin{subarray}{c}\|v\|\neq 0\\ v^{\top}\mathbf{1}=0\end{subarray}}\frac{v^{\top}L_{G}\,v}{v^{\top}L_{H}\,v}\;\leq\;\min_{\begin{subarray}{c}\|v\|\neq 0\\ v(i)\in\{0,1\}\end{subarray}}\frac{v^{\top}L_{G}\,v}{v^{\top}L_{H}\,v}.$ (9)
Restricting $v$ to $\{0,1\}$-valued entries treats $v$ as a cut and sums up the boundary edges in $G$ vs. $H$. In particular, if there exists a vertex $u$ of degree 1 in $G$ and the same edge connecting $u$ to its neighbor is preserved (with the same weight) in $H$, then assigning $v(u)=1$ and $v(i)=0$ for $i\neq u$ yields
$\frac{v^{\top}L_{G}\,v}{v^{\top}L_{H}\,v}\;=\;\frac{\omega_{u,\text{neighbor}}}{\omega_{u,\text{neighbor}}}\;=\;1,$ (10)
establishing $\lambda_{\min}(L_{G},L_{H})\leq 1$. Since a leaf vertex cannot become isolated in $H$ without invalidating the sparsifier, we also have $\lambda_{\min}(L_{G},L_{H})\geq 1$. Hence $\lambda_{\min}(L_{G},L_{H})=1$. If $L_{G}L_{H}^{+}$ thus has its smallest eigenvalue equal to 1, then $\lambda_{\max}(L_{G}L_{H}^{+})$ is precisely the relative condition number $\kappa(L_{G},L_{H})$. Consequently,
$\omega_{p,q}\,R_{p,q,H}\;\leq\;\frac{R_{p,q,H}}{R_{p,q,G}}\;\leq\;\kappa(L_{G},L_{H}),\quad\forall\,p,q\in V.$ (11)
In other words, no single edge’s effective-resistance ratio can exceed $\kappa(L_{G},L_{H})$, bounding the spectral distortion by the condition number.
Equation (5) indicates that an edge $e_{p,q}$ with a large product $w_{p,q}\,R_{p,q}$ can impose significant spectral distortion $\Delta_{K}$. Yet, directly computing the effective resistance $R_{p,q}$ for every inserted edge is prohibitively expensive on large graphs. To circumvent this bottleneck, we adopt a non-backtracking random walk approach that efficiently approximates or upper-bounds $R_{p,q}$ to enable rapid edge filtering.
(b) Non-Backtracking Random Walks and Approximate Resistance Distance: At the core of our method lies the observation that the accumulated effective resistance along the shortest path between two vertices $p$ and $q$ provides an upper bound on the true effective resistance $R_{p,q}$. Specifically, if one could identify the shortest path $P^{*}$ from $p$ to $q$, then the sum $\sum_{e\in P^{*}}r_{e}$ (where $r_{e}=1/w_{e}$) would serve as a valid upper bound on $R_{p,q}$, as resistance in parallel can only reduce total resistance.
However, directly computing shortest paths or all-pairs effective resistance remains computationally prohibitive in large-scale graphs. To circumvent this, we adopt a non-backtracking random walk (NBRW) approach, which probabilistically explores paths from $p$ without immediately revisiting the previous node, emulating a depth-first search pattern but in a randomized and lightweight manner. On tree-structured graphs, the path generated by NBRW from $p$ to $q$ corresponds exactly to the unique shortest path, making the accumulated resistance along the walk equal to $R_{p,q}$.
In general graphs, especially those sparsified using low-stretch spanning trees with off-tree edges, the NBRW path may not be the shortest. Nonetheless, the resistance accumulated along the walk still serves as a valid upper bound on the effective resistance, and—crucially—shorter paths are more likely to be sampled by repeated NBRWs. Thus, by initiating multiple non-backtracking walks from $p$, we estimate the resistance via:
$\tilde{R}_{p,q}:=\min_{1\leq i\leq s}\left(\sum_{e\in P^{(i)}}r_{e}\right)$ (12)
where each $P^{(i)}$ is a sampled non-backtracking path reaching $q$ and $s$ is the number of walkers.
This Monte Carlo strategy enables a practical, localized approximation of the effective resistance. The approximate spectral distortion for edge $e_{p,q}$ is then defined as $\tilde{\Delta}=\omega_{p,q}\tilde{R}_{p,q}$, and a distortion threshold $K$ is imposed to guide the random walk: walkers terminate their search once the cumulative approximate distortion exceeds this threshold ($\omega_{p,q}\tilde{R}_{\text{cumulative}}>K$), dynamically adjusting their search range rather than using a fixed step limit. In other words, the threshold $K$ acts as an effective proxy to filter edges in the sparsification process. For a newly introduced edge $e_{p,q}$:
  • If vertex $q$ is not reachable from vertex $p$ within the approximate distortion threshold $K$, the edge $e_{p,q}$ is considered spectrally critical and should be included in the sparsifier.
  • If vertex $q$ is reachable within the approximate distortion threshold $K$, the edge $e_{p,q}$ can be safely pruned without significantly impacting spectral similarity.
Moreover, the threshold $K$ provides a tunable tradeoff between sparsifier density and spectral accuracy:
  • Smaller $K$ results in more edges being retained, lowering the condition number but increasing sparsifier density;
  • Larger $K$ leads to more aggressive pruning, reducing density but potentially increasing the condition number.
To account for randomness and prevent false negatives (i.e., failing to reach $q$ despite it being within range), dyGRASS deploys $s$ independent walkers from $p$. The probability of all walkers missing $q$ decreases exponentially with $s$.
While the distortion threshold $K$ controls search adaptively, it does not bound the maximum number of steps. In regions where edge weights are large (i.e., resistances are small), many steps may be required before the accumulated distortion exceeds $K$. In fact, the worst-case number of steps can be loosely bounded by: $\frac{K}{\omega_{\min}/\omega_{\max}}$, where $\omega_{\min}$ is the minimum weight among candidate inserted edges, and $\omega_{\max}$ is the maximum edge weight in the graph. This implies that, even for moderate $K$, a walk might traverse a large portion of the graph in low-resistance regions.
To avoid such unbounded exploration and ensure computational efficiency, we introduce a fixed maximum step limit $T$ for each walker. Thus, a walker in dyGRASS may terminate under three conditions:
  1. 1.
    The cumulative distortion $\omega_{p,q}\tilde{R}_{\text{cumulative}}$ exceeds $K$,
  2. 2.
    The number of steps reaches the maximum threshold $T$,
  3. 3.
    The walker enters a leaf node (i.e., has no further neighbors to explore).
Importantly, when a walker terminates due to step limit $T$ without reaching $q$, dyGRASS conservatively assumes that $q$ lies outside the search region, and the edge $e_{p,q}$ is retained in the sparsifier. Although this may slightly increase sparsifier density, it ensures that no spectrally critical edges are erroneously removed—thus preserving or improving the spectral condition number.
$T$ effectively acts as a hyper-parameter in dyGRASS; it does not scale with graph size. Instead, $T$ controls how long each non-backtracking walker explores before terminating. A larger $T$ can accommodate a higher distortion threshold $K$, promoting more aggressive edge pruning (and potentially yielding a sparser graph) but risking an increase in the condition number. Conversely, a smaller $T$ strictly caps random-walk lengths and helps constrain the condition number to remain low, though it may retain additional edges.
The step limit $T$ caps the computational cost of each walk to at most $T$ steps, and with $s$ walkers per edge update, the total work is bounded by $s\times T$, which is a constant and independent of the graph size. This guarantees that dyGRASS achieves constant-time complexity $\mathcal{O}(1)$ per edge update in practice.
In our experiments, we fix $T=100$ as a reasonable default. This choice reflects practical observations that extremely high condition numbers are uncommon in many real-world EDA and network graphs, so a moderate $T$ sufficiently captures local structure information without excessive search overhead. Users can adjust $T$ based on their desired tradeoff between pruning aggressiveness (governed by $K$) and runtime constraints, all while preserving the $\mathcal{O}(1)$ update complexity characteristic of dyGRASS.
By combining non-backtracking random walks, a distortion threshold $K$, and a capped step limit $T$, dyGRASS achieves a practical approximation of effective resistance distances in large graphs. This framework avoids the computational cost of exact shortest-path or pseudo-inverse methods, preserving high spectral fidelity at minimal overhead.
III-C Decremental spectral graph sparsification
Consider an edge $e_{p,q}\in E$ to be deleted from the original graph $G^{(0)}$. The goal is to efficiently update $H^{(0)}$ while preserving the spectral similarity between $G$ and $H$. Our decremental method will first determine if $e_{p,q}$ belongs to the $H^{(0)}=(V,E^{\prime(0)},w^{\prime(0)})$:
if $\mathbf{e_{p,q}\notin E^{\prime(0)}}$: $H^{(0)}$ remains unchanged, since $e_{p,q}$ is not part of the sparsifier.
if $\mathbf{e_{p,q}\in E^{\prime(0)}}$: The sparsifier must be updated by recovering critical edge(s) from $G^{(1)}$ to $H^{(0)}$, so as to preserve crucial distances in the new graph.
Distance estimation using non-backtracking random walks: Let us assume that $e_{p,q}\in E$ is deleted from $G^{(0)}$ and also exists in the spectral sparsifier $H^{(0)}$. Simply removing this edge from $H^{(0)}$ may isolate certain vertices or significantly perturb their distances. In other words, deleting a sparsifier edge can disrupt connectivity and inflate effective resistances between previously connected vertices.
To address these challenges, it is necessary to recover edges from the updated original graph $G^{(1)}$ (after the deletion) and add them back into $H^{(0)}$. We achieve this by performing localized NBRWs:
  • Multiple Walkers with Step Limit $T$: For the deleted edge $e_{p,q}$, we launch several non-backtracking walkers from $p$ (or from $q$). Each walker proceeds up to $T$ steps without traversing the same edge in consecutive steps.
  • Accumulated Resistance: As a walker moves along a path $P$, it accumulates the sum of resistances ($1/w_{e}$) for the edges $e\in P$. This sum represents a distance measure in the updated graph $G^{(1)}$.
  • Selecting the Minimum-Resistance Path: Among all paths that successfully reach $q$, only the path with the smallest total resistance is chosen for recovery. The edges along this path are added to the sparsifier to restore the updated resistance distances between $p$ and $q$.
In the event that no walker successfully reaches $q$ within $T$ steps, it implies that the resistance distance between $p$ and $q$ after edge deletion is sufficiently large and no specific edge recovery is required to maintain spectral fidelity. dyGRASS still recovers a few local edges near $p$ and $q$ to prevent the creation of isolated components from forming in the sparsifier. This approach ensures the updated sparsifier reflects the structural changes in $G^{(1)}$, preserving spectral integrity without exhaustive searches. This localized NBRW-based scheme ensures the structural and spectral integrity of the updated sparsifier when handling edge deletions. By bounding each walker’s path length to $T$, the time spent exploring is limited, yielding an efficient update procedure that seamlessly complements dyGRASS’s incremental phase.
III-D The dyGRASS algorithm flow
Algorithm 1 dyGRASS: Dynamic Spectral Sparsification
1:  Input: $G^{(0)}$: Original graph, $H^{(0)}$: Sparsifier, EdgeStream: Stream of updates, distortion threshold $K$, maximum step size $T$, number of walkers $s$, update type $choice$
2:  Output: $H^{(t)}$: Updated sparsifier
3:  Initialize $H^{(t)}\leftarrow H^{(0)}$
4:  for each update $(\text{type},(u,v))$ in EdgeStream do
5:     if choice == “insertion” then
6:        1. Add $(u,v)$ to $G^{(t)}\to G^{(t+1)}$
7:        2. Perform non-backtracking random walks from $u$ to $v$ on $H^{(t)}$, using $K$, $T$, and $s$
8:        if $v$ not reachable by any walker then
9:           then add $(u,v)$ to $H^{(t)}$
10:        end if
11:     else if choice == “deletion” then
12:        Remove $(u,v)$ from $G^{(t)}\to G^{(t+1)}$
13:        if $(u,v)\in H^{(t)}$ then
14:           1. Remove $(u,v)$ from $H^{(t)}$
15:           2. Perform non-backtracking random walks between $u$ and $v$ on $G^{(t+1)}$, up to step limit $T$
16:           3. Select a path with minimum accumulated resistance in $G^{(t+1)}$
17:           4. Recover edges from this path into $H^{(t)}$
18:        end if
19:     end if
20:     Update $H^{(t)}\to H^{(t+1)}$
21:  end for
22:  return  $H^{(t)}$
The flow of the proposed dynamic graph sparsification algorithm (dyGRASS) is shown in Algorithm 1 and described as follows:
Lines 1–3: The inputs include the original graph $G^{(0)}$ and its initial spectral sparsifier $H^{(0)}$ precomputed using GRASS method [12], and a stream of edge updates (EdgeStream). The threshold $K$ and step limit $T$ define local exploration for each walker, and $s$ denotes the number of walkers.
Lines 5–9: For edge insertions, dyGRASS adds $e_{u,v}$ to $G^{(t)}$ and checks via NBRW in $H^{(t)}$. if $(u,v)$ is unreachable within $K$ or step $T$, it is considered critical and is kept; otherwise, it can be pruned.
Lines 12–17: For a deletion, dyGRASS removes $e_{u,v}$ from $G^{(t)}$. if $e_{u,v}$ is also in $H^{(t)}$, it is removed and a minimum resistance path is recovered via random walks in $G^{(t+1)}$.. If no path is found, adjacent edges near $(u,v)$ will be recovered to prevent forming isolated components.
The per-update time complexity of dyGRASS is effective $O(1)$, under fixed parameters $T$ and $s$, independent of graph size.
III-E A Case Study for Dynamic Spectral Graph Sparsification
Figure 3: Schematic of the dyGRASS algorithm illustrating incremental updates of $e_{25,17}$ and $e_{25,9}$, as well as decremental updates of $e_{16,17}$ and $e_{24,25}$.
Fig. 3 presents an example of dyGRASS managing both incremental and decremental updates on a simple undirected graph. In Fig. 3(a), two edges, $(25,17)$ and $(25,9)$, are introduced in the original graph $G^{(0)}$, producing the updated graph $G^{(1)}$. Non-backtracking random walkers begin at vertex 25 in the current sparsifier $H^{(0)}$, using a distortion threshold $K=4$ (equivalent to a maximum walk length of four steps), as illustrated by the blue dashed circle.
From these walks, vertex 9 remains unreachable, indicating that $(25,9)$ is spectrally critical that must be included in the updated sparsifier. Conversely, since vertex 17 is reached by at least one walker, the newly introduced edge $(25,17)$ is excluded from the sparsifier. The resulting sparsifier is denoted as $H^{(1)}$.
In Fig. 3(b), two edges, $(16,17)$ and $(24,25)$, are deleted from $G^{(0)}$, producing another updated graph $G^{(1)}$. The sparsifier after edge deletion but before recovery is shown as $H^{(0)}$. Since $(24,25)$ is not in $H^{(0)}$, no further action is required. However, the deleted edge $(16,17)$ is in $H^{(0)}$, prompting a sparsifier update. Walkers are thus initialized from vertex 16 in $G^{(1)}$, each limited to four steps. Among the discovered paths, the one with the minimum accumulated resistance distance includes $(16,21)$, $(21,22)$, and $(22,17)$ (highlighted in green). Only $(16,21)$ is not already in $H^{(0)}$ and is therefore recovered into $H$. The updated sparsifier $H^{(1)}$ is then obtained.
This example illustrates how dyGRASS dynamically identifies and preserves spectrally significant edges under incremental and decremental operations, maintaining strong spectral similarity with the evolving original graph.
III-F GPU-accelerated random walks
Although dyGRASS maintains an $O(1)$ update time per edge in theory, the actual constant factors can become significant if the step limit $T$ or the number of walkers $s$ grows large. For incremental updates, a high condition number may necessitate an increased $T$ to correctly capture high-distortion edges. In decremental updates, despite the original graph often having a smaller radius, deploying additional walkers can still increase the total workload. Consequently, each update can involve a substantial amount of random-walk exploration when $T$ or $s$ is large.
While multi-threading on CPUs can initiate multiple walkers in parallel, it may not scale well for large $s$ or large batches of updates. In contrast, the SIMT (Single Instruction, Multiple Threads) architecture of modern GPUs is inherently well-suited to random-walk parallelization [37]. By assigning each GPU thread a walker, we can process many edges (or many walks) in parallel, significantly reducing the wall clock time for each batch of updates.
However, supporting dynamic updates efficiently on the GPU presents two key challenges: (1) fast neighbor queries during the walk, and (2) fast edge insertions/deletions for both incremental and decremental updates. Traditional data structures like the Compressed Sparse Row (CSR) format enable efficient neighbor access in $O(1)$ time but require costly full reconstruction upon edge insertions or deletions. In contrast, the Coordinate (COO) format supports more flexible edge updates but suffers from inefficient neighbor queries due to its unsorted and unindexed layout. To address these issues, recent GPU frameworks have proposed more general-purpose dynamic graph data structures optimized for high-throughput workloads. For example, LPMA [38] improves update performance through a leveled tree layout and localized re-balancing strategies, while a hash-based design [39] assigns each vertex a dedicated hash table to enable constant-time edge insertions and deletions. These systems achieve impressive scalability and support large-scale dynamic workloads efficiently. However, these solutions are primarily optimized for high-churn settings where frequent large batches of updates dominate the workload. In our application, edge updates are relatively moderate—typically around 25% of the vertex count—and query performance is more critical than raw update speed.
Figure 4: Schematic of dynamic CSR data structure illustrating insertion of $(u,v_{4})$ and deletion of $(p,q_{2})$.
To this end, we design a lightweight, self-defined dynamic CSR structure that replaces traditional row offsets with a vertex pointer list that directly references each vertex’s adjacency lists in global memory. During initialization, we allocate an extra memory block to accommodate future edge insertions. As shown in Fig. 4, incremental updates write new neighbor sub-arrays into this extra space and update the pointers in the vertex pointer array accordingly. Decremental updates, by contrast, can be handled in-place without additional memory overhead. This design preserves efficient neighbor lookups while permitting updates without reconstructing the entire data structure. In the GPU setting, we adopt a batch-processing model to maximize throughput and reduce host-device synchronization. Unlike the CPU version of dyGRASS, which applies updates immediately, the GPU version delays updates until the end of a batch. This deferral may slightly affect sparsifier quality in the short term, but it substantially improves overall efficiency by minimizing PCIe traffic and increasing GPU utilization. To parallelize random walks, we assign one GPU thread per walker. Each thread performs a non-backtracking random walk independently, using the dynamically updated adjacency array. Within a thread block, all walkers handle the same edge insertion or deletion in parallel. At each step, a thread reads the degree of its current vertex, follows the corresponding pointer in the vertex pointer array to retrieve the neighbor sub-array, and selects a neighbor other than the one visited in the previous step—thereby enforcing the non-backtracking constraint. This lightweight approach balances dynamic updates with fast neighborhood exploration, making it a practical choice for GPU-accelerated random walks in dyGRASS.
IV Experimental results
This section outlines the findings from a series of experiments designed to evaluate the effectiveness and efficiency of the proposed dyGRASS. These experiments are conducted using a variety of public domain sparse matrices 111available at https://sparse.tamu.edu/ originating from various real-world applications, such as circuit simulation and finite element analysis. All experiments were conducted on a Linux Ubuntu $24.04$ system with $64$ GB of RAM, a $3.4$ GHz $8$-core AMD R7 5800X3D CPU, and an Nvidia RTX 4090 GPU. The state-of-the-art spectral sparsification tool GRASS [12]  [^2] https://sites.google.com/mtu.edu/zhuofeng-graphspar/home has been used to generate the initial sparsifier, while the state-of-the-art incremental spectral sparsification tool inGRASS [32] served as the baseline for assessing the performance and scalability of dyGRASS.
TABLE II: Graph density and spectral similarity of the initial (updated) spectral sparsifiers $H^{(0)}$ ($H^{(10)}$).
Test Cases $|V|$ $|E|$ $D_{H}^{(0)}\rightarrow D_{H}^{(10)}$
$\kappa(L_{G}^{(0)},L_{H}^{(0)})$
$\rightarrow$
$\kappa(L_{G}^{(10)},L_{H}^{(0)})$
[…]
TABLE III: Comparison of Dynamic Graph Spectral Sparsification Outcomes through $10$-Iterative Updates applying iinGRASS and dyGRASS.
inGRASS.incremental() [32] dyGRASS.incremental() dyGRASS.decremental()
Test Cases $T_{setup}$ $T_{update}$ $D_{H}^{(10)}$ $\kappa(L_{G}^{(10)},L_{H}^{(10)})$ $T_{update}$ $D_{H}^{(10)}$ $\kappa(L_{G}^{(10)},L_{H}^{(10)})$ $T_{update}$ $D_{H}^{(20)}$ $\kappa(L_{G}^{(20)},L_{H}^{(20)})$
G2_circuit 1.77s 0.01s 13.97% 95.41 0.019s 12.10% 71.91 0.022s 12.87% 70.15
G3_circuit 23.5s 1.20s 13.91% 88.35 0.088s 13.58% 88.39 0.155s 14.37% 113.51
fe_4elt 0.12s 0.01s 15.01% 136.32 0.010s 13.87% 98.59 0.011s 14.40% 98.54
fe_ocean 1.77s 0.11s 16.34% 206.97 0.040s 13.39% 209.70 0.026s 14.29% 194.92
fe_sphere 0.16s 0.01s 14.07% 173.35 0.010s 13.88% 144.77 0.011s 14.41% 144.68
delaunay_n18 3.57s 0.20s 15.29% 171.04 0.024s 14.56% 132.77 0.038s 15.05% 132.70
delaunay_n19 7.79s 0.39s 18.73% 138.18 0.040s 15.44% 127.23 0.069s 15.92% 129.07
delaunay_n20 17.02s 0.75s 18.43% 148.13 0.075s 15.09% 146.61 0.127s 15.58% 146.60
delaunay_n21 36.69s 1.60s 18.00% 159.90 0.145s 14.50% 147.54 0.273s 14.99% 147.10
delaunay_n22 80.26s 4.32s 18.19% 176.14 0.283s 14.40% 164.36 0.497s 14.90% 164.37
M6 66.53s 3.72s 17.73% 232.50 0.235s 13.43% 173.64 0.474s 13.92% 173.63
 [^3] SP 67.00s 3.68s 14.13% 232.57 0.251s 13.61% 182.46 0.475s 14.11% 183.39
AS365 73.26s 4.30s 14.63% 196.76 0.258s 13.6% 183.64 0.490s 14.10% 183.33
NACA15 17.14s 1.00s 15.19% 219.66 0.074s 13.94% 150.16 0.137s 14.44% 150.17
IV-A Dynamic spectral sparsification using dyGRASS
We evaluate the performance of dyGRASS by comparing the off-tree edge density of the sparsifier, which is defined as $D_{H}:=\frac{|E|}{|V|}-1$ the same relative condition number is achieved using inGRASS and dyGRASS methods. The relative condition number $\kappa(L_{G},L_{H})$ can be deployed to quantify the spectral similarity between the original graph $G$ and its graph sparsifier $H$ [11, 12]. A smaller relative condition number implies higher spectral similarity between the graph $G$ and its sparsifier $H$.
As shown in Table II, the graph sparsifier density increases significantly ($D_{H}^{(0)}\rightarrow$ $D_{H}^{(10)}$) when all the edges are inserted into $H^{(0)}$ through $10$-iterative updates without an edge filtering mechanism. On the other hand, if we keep the initial graph sparsifier $H^{(0)}$ unchanged through $10$-iterative updates, the relative condition number $\kappa(L_{G}^{(10)},L_{H}^{(0)})$ will grow significantly from the initial condition number $\kappa(L_{G}^{(0)},L_{H}^{(0)})$, implying increasing mismatches between graph $G^{(10)}$ and the initial sparsifier $H^{(0)}$.
Table III compares GPU-accelerated dyGRASS and inGRASS [32] over $10$ incremental updates. During the ten incremental updates, approximately $25\sim 30\%$ edges in total are inserted into $G^{(0)}$, as shown in column $4$ of Table II. Subsequently, we also perform ten decremental updates using dyGRASS, in which $1\%$ edges in total are randomly removed from $G^{(10)}$. Each incremental or decremental update is configured to produce sparsifiers matching the original spectral similarity $\kappa(L_{G}^{(0)},L_{H}^{(0)})$. As shown in Table III, dyGRASS.incremental() consistently achieves lower or similar sparsifier densities compared to ones produced by inGRASS.incremental(). Additionally, for the subsequent ten decremental updates, dyGRASS.decremental() is always able to efficiently preserve low densities of the updated spectral sparsifiers $(D_{H}^{(20)})$ while preserving high spectral quality.
IV-B Sparsifier densities after multiple incremental updates
Figure 5: Comparison of sparsifier densities in 10 incremental updates
Fig. 5 illustrates dyGRASS’s dynamic updating capability in ten iterations for handling the circuit graph ”G2_circuit”. Since dyGRASS can capture the latest structural changes (e.g., decreasing resistance diameters) in every update iteration via localized random walks, it allows the newly inserted edges to be selected more optimally considering the most recent updates. As a result, dyGRASS is able to achieve much lower sparsifier density than inGRASS as more updates are made. In contrast, when more edges are inserted into the graph, the resistance distances between nodes will monotonically decrease, leading to an increasing overestimation of the spectral distortion of each new edge since inGRASS keeps processing newly inserted edges using the dated resistance embedding.
TABLE IV: Solver Iterations and Sparsifier Densities with/without dyGRASS Updates.
Initial sparsifier No update dyGRASS update
$\kappa(L_{G}^{(0)},L_{H}^{(0)})$ $D_{H}^{(0)}$Iters $\Delta D_{G}^{(10)}$Iters $D_{H}^{(10)}$Iters
107.69 13.77% 94 24.10% 273 17.4% 103
183.18 8.52% 122 24.10% 336 12.33% 120
407.92 4.92% 171 24.10% 426 9.64% 152
IV-C Preconditioned Solver Iterations
To solve a linear system $L_{G}\,x=b$ more efficiently, we use a spectral sparsifier $H$ as a preconditioner. That is, we replace $L_{G}$ by $L_{H}$ in each iterative step, effectively solving $L_{H}^{-1}L_{G}\,x=L_{H}^{-1}b$. The convergence rate hinges on the condition number $\kappa(L_{G},L_{H})$, so a high-quality sparsifier leads to fewer solver iterations. In this experiemnt, the PCG solver terminates when the relative residual norm $<10^{-8}$.
Table IV considers three initial sparsifiers of the “AS365” graph, each with a different condition number and corresponding iteration count (#Iters). After inserting $24.1\%$ new edges into $G^{(0)}$ to form $G^{(10)}$, leaving the preconditioner $H^{(0)}$ unchanged (No update) inflates the iteration count significantly. In contrast, applying dyGRASS updates (dyGRASS update) slightly increases the off-tree density but drastically reduces the iteration count. This demonstrates how dyGRASS retains the crucial edges for preconditioning while pruning less important ones, maintaining a low-density sparsifier that preserves robust solver performance despite added edges.
IV-D Runtime scalability of dyGRASS
Figure 6: Runtime scalability comparison between inGRASS and dyGRASS.
In Fig. 6, we compare the runtime scalability of dyGRASS and inGRASS across various real-world graphs. Over ten iterative incremental updates, the GPU implementation of dyGRASS achieves approximately a 70$\times$ speedup over its CPU counterpart. Compared to inGRASS, dyGRASS offers about a 10$\times$ improvement for large graphs. Although dyGRASS maintains $O(1)$ time complexity per edge update, the total runtime increases with graph size due to the growing number of updates. Specifically, we fix the update density at approximately 25% of the number of vertices. As a result, larger graphs involve proportionally more edge updates per iteration, leading to an increase in total update time.
V Conclusion
This work presents dyGRASS, an efficient and scalable algorithm for dynamic spectral sparsification of large undirected graphs. By employing random-walk-based techniques for incremental and decremental updates, dyGRASS can efficiently identify spectrally critical edges essential for preserving graph spectral properties while pruning redundant edges to maintain desired sparsifier density. The proposed dyGRASS algorithm also leverages GPU acceleration to efficiently process a large number of dynamic updates. Our extensive experimental results demonstrate that dyGRASS outperforms state-of-the-art methods, inGRASS, in both runtime and solution quality. It achieves significant speedups while dynamically maintaining spectral fidelity across various graph types, including those from circuit simulations, finite element analysis, and social networks. Overall, the dyGRASS framework bridges the gap between theoretical advancements in spectral graph theory and practical EDA applications, providing a robust and scalable solution for compressing dynamically evolving graphs.